OpenStudy (whynot):

integrate x^3 dx @ integrate x^3 dx @ @Mathematics

OpenStudy (anonymous):

use the power rule backwards

OpenStudy (anonymous):

\[ \frac{x^4}{4} +C \]

OpenStudy (whynot):

x^4/4?

OpenStudy (anonymous):

\[\frac{d}{dx}[x^n]=nx^{n-1}\] \[\int x^n dx=\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

yes you have it

OpenStudy (whynot):

thanks a bunch:)